OpenStudy (anonymous):

function is: 9x^3+10x^2+5 find the equation of the tangent line at x=0, please show steps

OpenStudy (campbell_st):

can you find the 1st derivative...?

OpenStudy (anonymous):

Yup

OpenStudy (campbell_st):

ok... so what did you get...?

OpenStudy (anonymous):

the derivative stuff isn't the problem. I'm only slightly confused on a certain part

OpenStudy (campbell_st):

the 1st derivative gives the equation for the slope of the tangent...

OpenStudy (anonymous):

Yes, yes I understand that

OpenStudy (campbell_st):

ok... so you have \[\frac{dy}{dx} = 27x^2 + 20x\]

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

ok... so find the slope of the tangent at x = 0 substitute x = 0 into the 1st derivative... so m = 0 you are dealing with a horizontal line does that make sense..?

OpenStudy (anonymous):

Yes, I've gotten all that

OpenStudy (anonymous):

I just wasn't sure if that was correct. It just didn't look right

OpenStudy (campbell_st):

ok.. you need to y value for the point where x = 0 so substitute x = 0 into the original equation y = 9*0^3 + 10*0^2 + 5 so y = 5 the point you are working with is (0, 5) does that make sense..?

OpenStudy (anonymous):

but since you confirmed my answer, I guess there's not a real problem

OpenStudy (anonymous):

Yep, I have all that :)

OpenStudy (campbell_st):

well the slope intercept form is y = mx + b in this question y = b substitute x = 0 and y = 5 gives b = 5 so the line is y = 5

OpenStudy (anonymous):

Ok good good yes yes I have thees :D

OpenStudy (anonymous):

w00t

OpenStudy (campbell_st):

well done...

OpenStudy (anonymous):

wouldn't it be -5 (unless I'm crazy)

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

I goofed

OpenStudy (campbell_st):

I think its y = 5 are you happy with the point being (0, 5)...? and m = 0 so y = 0x + b or 5 = 0*0 + b b = 5

OpenStudy (anonymous):

It's just five haha...

OpenStudy (anonymous):

And yes, I'm happy with it. I needed to make sure that I wasn't going completely insane. Thank you :)

OpenStudy (campbell_st):

glad to help